Unit 22 Homework Assignment  Physics141-Summer 2009

Unit 22 homework after session 1

• Read chapter 26 in the text book Understanding Physics up to section 26-5.

• Complete unit 22 homework #1: Batteries, Bulbs and Current. Unit 22 homework after session 2

• Read the rest of chapter 26 in the text book Understanding Physics.

• Complete unit 22 homework #2: Direct Circuits.

• Work textbook problems 26-19, 26-22 and 26-43.

All Unit 22 homework problems and the unit 22 activity guide are due on the first day of unit 23.